Suppose Japan has a GDP of $5 trillion, and that its national savings rate is 25%. Assuming Japan is an open economy, i. Calculate Japan’s investment if net exports are 1% of GDP ii. Calculate Japan’s exports if imports are valued at $650 billion
Answer #1

a)National saving = 5 trillion*25% = 1250 billion

Net exports = 1% of GDP

so, NX= 5 trillion * 1% = 50 billion

Since Y=C+I+G+NX

Y-C-G=I+NX

1250=I+50

I=1250-50 = $1200 billion or $1.2 trillion

b) Imports = 650 billion

Since NX = Exports - Imports

So, 50 billion = E-650 billion

Exports = 50+650 = $700 billion